OpenStudy (anonymous):

Hey there, Im working a problem finding the dy/dx for y=(3x-2)/(x^2+2x-3). I did the quotient rule and I'm stuck after simplifying the top.)

zepdrix (zepdrix):

\[\large y=\frac{3x-2}{x^2+2x-3}\] \[\large y'=\frac{\color{royalblue}{(3x-2)'}(x^2+2x-3)-(3x-2)\color{royalblue}{(x^2+2x-3)'}}{(x^2+2x-3)^2}\] \[\large y'=\frac{\color{royalblue}{(3)}(x^2+2x-3)-(3x-2)\color{royalblue}{(2x+2)}}{(x^2+2x-3)^2}\] So you're having trouble simplifying it from here?